Hello! I have a word problem that I am having trouble with on my algebra homework... If you would help that would be awesome!
Here it is:
Ben adds a 65% fruit juice solution to 15mL of a drink that is 10% fruit juice. How much of the 65% fruit juice solution must be added to make a punch that is 35% fruit juice?
Thank you!

Answer by mananth(16949)

--- percent ---------------- Amount
Juice A 10 ---------------- 15 ml
Juice B 65 ---------------- x ml
Mixture 35 ---------------- 15 + x ml

10 * 15 + 65 x = 35 ( 15 + x )
150 + 65 x = 525 + 35 x
65 x -35 x = -150 + 525
30 x = 375
/ 30
x = 12.5 ml Juice B

m.ananth@hotmail.ca